OpenStudy (anonymous):

Find the derivative of w(q) = 7ab6q I have tried 42b^5q and 6b^5q and neither have worked. Please help!

OpenStudy (anonymous):

So I found the answer, it was 7ab^6. But can someone please help me figure out why it is the answer!

OpenStudy (anonymous):

is it 7ab raised to 6 ??

OpenStudy (anonymous):

well if it so... u should knw that the derivative of a first degree is 1.... and if it has a coefficient then it implies itz derivative is coefficient is multiplied by one nw try thinkin over it

OpenStudy (cwrw238):

the answer is 7ab^6 as its the coefficient of the variable q and q is of degree 1

OpenStudy (cwrw238):

compare for example 7q - the derivative is 7

OpenStudy (anonymous):

right, I know all of that, and thanks for the responses!! I should have been clearer, why doesn't 7a become 7, and b^6 become 6b^5 for an overall answer of 42b^5.

OpenStudy (anonymous):

itz like... a and b are constants which implies 7ab^6 is itself a constant...simple

OpenStudy (cwrw238):

the question asks for a derivative with respect to q - we can take the 7ab^2 to be constants

OpenStudy (anonymous):

what u shud understand first is if the function is f(x)... x is the variable and in W(q) q is the variable/.... and generally the alphabets a and b denotes constants

OpenStudy (anonymous):

Ok, thanks guys!! It makes sense now...I was so confused lol

OpenStudy (aum):

\[ \large w(q) = (7ab^6) * q \\ \large \frac{dw(q)}{dq} = \frac{d}{dq}((7ab^6) * q) = 7ab^6 * \frac{d}{dq}q = 7ab^6 * 1 = 7ab^6 \]